# InComeS: Integrating Compression and Selection Mechanisms into LLMs for Efficient Model Editing

### v2

@@ 4.3 Ablation study & Analysis
Method Model DUNE MQuAKE New info Scientific R. Debiasing 2-edits 3-edits 4-edits InComeS Llama-3.2-1B 57.76 56.46 46.14 53.93 52.79 52.73 - w/o zero-gist 56.85 57.46 45.72 47.48 41.71 32.57 - w/ full model 56.25 54.28 38.31 51.91 52.59 52.53 - w/o loss on query 58.89 57.49 46.25 49.63 46.44 38.28 - w/ golden loss 56.46 53.62 43.05 55.39 50.24 40.06
Table 4: Ablation and Analysis experiments, the edit batch size is 100 for all results.
Inclusion of zero gist
The motivation of including the zero-gist mechanism is to ensure that context-independent tokens can bypass the influence of the edit contexts. To assess the impact of zero-gist, we train a model without this mechanism and evaluate it on DUNE [1] and MQuAKE [49] (see the “w/o zero-gist” line in Table 4). The results show a notable performance drop, particularly in multi-hop scenarios, suggesting that the cross-attention calculations may sometimes interfere with ordinary generation and our zero-gist strategy can mitigate this issue by allowing tokens to “attend to nothing”.
@@
We present the reason for our decision to use the KV cache from the second half of the model layers. To investigate this, we train a model using the KV cache from all layers and evaluate it on 1000 instances from ZsRE [20]. We record the probabilities allocated to the zero-gist in the cross-attention modules, as shown in Fig. 4b. The result shows that the zero-gist probabilities in layers 7-15 are generally lower than those in layers 1-6, and there is a notable drop in the zero-gist probability at layer 7. This suggests that, even when trained to use the full-model KV cache, the model mainly relies on information from deeper layers, since higher probabilities of the zero-gists indicate lower utilization of the actual edit contexts. A possible explanation is that more information is accumulated in the deeper layers, which aids both compression and selection processes. To verify our analysis, we also test the full layer trained model (see the “w/ full model” line in Table 4) under a batch editing setting, and observe a general decrease across all metrics. Additionally, restricting the KV cache to only the second half of the model could provide efficiency benefits with lower memory and computation costs.
Applying loss on queries
@@
Applying a small temperature to the gist cross-attention sharpens the probability distribution over the gist KV caches, which facilitates the model’s ability to retrieve the correct information. We determine the appropriate temperature based on entropy, which has been shown to be an important factor in attention mechanisms [47]. Specifically, we aim to keep the cross-attention entropy close to its optimal value, which occurs when it only needs to attend to one edit. To achieve this, we select 1000 instances from ZsRE [20] and calculate the entropy of the edit batch size 1. We then calculate the entropy for larger edit batch sizes (10, 20, 40, 60, 80, 100, 200, 300) and find the temperatures that align their entropy with the optimal case via gradient descent  [^6] Note that all entropy is calculated in a way that the query and answer part are just a copy of the context. We report the calculated temperature in Figure 4a. As expected, the temperature decreases as the edit batch size increases, but interestingly, it gradually converges to a specific value. Specifically, layers 9, 10, and 14 finally converge to around 0.5. To encourage more decisive selection, we slightly lower this value and set the temperature to $T=0.45$.
Information flow on tokens
@@
Side effect analysis
The side effect analysis is provided in Appendix D.2.
4.4 Efficiency Analysis
Finally, we present the efficiency analysis for our method. By default, the individual edit length is around 10 to 11. We first compare the efficiency of our method with the efficiency of other knowledge editing methods. Table 5 reports the time required to perform 100 edits for each method. Our method has significantly better efficiency than the other presented editing methods. Additionally, compared to ICL, our approach only needs to maintain the KV cache of the gist representations from the deeper half layers, resulting in substantially lower memory cost. To verify our method’s superiority in efficiency on long context, we further conduct experiments on scaled context length (Table 6). The result demonstrates the efficiency advantage of our method in both the encoding and decoding stages. More detailed analysis can be found in Appendix D.3.
Model InComeS-Compression ICL-Prefilling FT-M LoRA [16] MEMIT [25] EMMET [11] Llama-3.2-1B 0.0326 0.8934 3.4124 33.5412 112.2238 158.6524 Qwen2.5-7B 0.1071 0.9082 28.6132 122.1876 423.4823 512.4235
Table 5: Measured time (seconds) for 100 edits.
Number of edits Model Encoding Decoding InComeS-Compression ICL-Prefilling InComeS-Selection ICL-Generation (with prefilled cache) 1k Llama-3.2-1B 0.2108 1.0413 0.0274 0.1555 2k 0.4051 1.2165 0.0297 0.3545
Table 6: Scaled efficiency comparison (seconds) between InComeS and ICL.
@@ 5 Related Work
The area of knowledge editing (or model editing) has experienced a thriving development in recent years. Researchers have explored various directions in this area. One typical direction is to adopt external memory for the edits. The memory formats applied by different researchers are diverse. Methods like SERAC [29], IKE [48], MeLLo [50] adopt explicit non-parametric memory, which stores specific edit instances, and a retriever that is responsible for recalling relevant edits from the memory. For example, IKE uses KNN, and SERAC applies a trained classifier. Another line of work applies implicit parametric memory. CaliNET [7], T-Patcher [17] embeds the knowledge into a fixed number of neurons and adds them to the model. GRACE [12] adopts a discrete key-value codebook with the value optimized for the desired knowledge. MELO [44] applies dynamic LoRA blocks and indexes them via an internal vector database. KE [4], MEND [28] train a separate meta-model for editing. Another popular direction is to merge knowledge into the model directly. Methods like KN [3], ROME [24], R-ROME [10], MEMIT [25], PMET [22], CoachHooK [21], and AlphaEdit [8] perform editing by tweaking the located FFN part of the model directly. However, some studies reveal that these methods could potentially bring about side effects in the original model [9, 33], leaving the real effectiveness of these methods to be further investigated.
@@ 6 Conclusion
In this paper, we propose InComeS, a scalable model editing method that integrates compression and selection mechanisms directly into the LLMs. InComeS adopts a context compression technique to condense the editing context to KV representations on top of the introduced gist tokens and takes advantage of the compressed KVs to efficiently retrieve the relevant editing context information. Experiments on four different and complex editing settings demonstrate the superiority of our method for comprehensively editing. Further Analysis and ablations validate each component of InComeS and demonstrate the advantage of our method in efficiency.
Appendix A Limitations
@@
Due to the limited computational resources, we only extend the model size to 7B and leave the larger model size to future work. We are aware that the original gisting work [30] conducts experiments on three model architectures, i.e., encoder-decoder, encoder-only, and decoder-only. In this work, we determine to focus on the decoder-only autoregressive architecture as it is the structure used by most of the popular models nowadays [42, 31, 5].
@@
In this work, we maintain the compression rate to roughly around 12:1 (one edit, which contains around 12 tokens for all testing datasets except DUNE [1], corresponds to one gist token), as one edit represents a fine-grained piece of information. However, we believe it is necessary to investigate the impact of lowering the compression rate, since it potentially helps extend the length of a single edit [6].
@@
Appendix B Training details
@@
InComeS is trained on around 1.5 billion tokens, which mainly come from summarization and QA datasets. Specifically, for summerization datasets, we select $4.5e^{6}$ instances from S2ORC [23], $1.15e^{6}$ instances from AG News Corpus  [^7] https://huggingface.co/datasets/sentence-transformers/agnews; and for QA datasets, we use squad [36], a modified version  [^8] https://huggingface.co/datasets/LLukas22/nq-simplified of the natural question dataset [19], OpenBookQA [27], QASC [18], MedMCQA [32], and NetEval [26]. We also include the training split of ZsRE [20], COUNTERFACT [25], $Wiki_{counterfact}$ from EasyEdit framework [39].
@@
We use a cosine linear-warmup scheduler for both models, with a maximum learning rate $1e^{-5}$ and a minimum learning rate $1e^{-6}$ for Llama-3.2-1B and a maximum $5e^{-6}$ and a minimum $1e^{-6}$ for Qwen2.5-7B. To improve the model’s robustness and sample them at a predefined rate during training, the batch size is dynamically sampled from a predefined set rather than a fixed number. Specifically, the predefined set for batch size is 8, 16, 32, 64, and 128, and their corresponding sample rates are 0.05, 0.05, 0.05, 0.15, and 0.7. We adopt DeepSpeed [35, 34, 37] and Liger Kernel [15] with 8 Nvidia GPUs for distributed training. Overall, the training takes around 11 hours for Llama-3.2-1B and 35 hours for Qwen2.5-7B.
Appendix C Experiment details
@@
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Locality Edit Success Portability Locality Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 - 30.06 / 30.06 40.17 / 40.17 - FT-M 97.02 / 94.58 53.43 / 47.51 49.01 / 20.54 99.81 / 95.94 62.80 / 54.84 75.59 / 59.28 LoRA [16] 98.91 / 82.61 52.87 / 43.84 23.31 / 14.94 99.86 / 93.18 57.43 / 44.85 34.94 / 37.41 ROME [24] 94.33 / 44-33 40.44 / 44-33 26.13 / 44-33 95.41 / 44-33 46.04 / 44-33 40.19 / 44-33 R-ROME [10] 94.31 / 44-33 40.65 / 44-33 25.83 / 44-33 95.29 / 44-33 46.46 / 44-33 39.95 / 44-33 MEMIT [25] 44-33 / 66.94 44-33 / 23.51 44-33 / 11.12 44-33 / 58.79 44-33 / 25.68 44-33 / 30.10 EMMET [11] 44-33 / 27.02 44-33 / 08.86 44-33 / 100.0 44-33 / 15.96 44-33 / 06.84 44-33 / 100.0 MEND [28] 44-33 / 26.66 44-33 / 21.06 43-33 / 13.54 34-33 / 43.33 34-33 / 30.77 34-33 / 38.75 GRACE [12] 33.27 / 25.06 14.33 / 10.51 28.71 / 11.43 32.00 / 24.44 12.73 / 10.91 24.12 / 10.12 KN [3] 20.60 / 34-33 17.16 / 34-33 19.46 / 43-33 16.01 / 43-33 06.70 / 43-33 21.23 / 34-33 IKE [48] 61.70 / 34-33 45.55 / 34-33 48.80 / 43-33 59.15 / 43-33 57.39 / 43-33 40.21 / 34-33 SERAC [29] 89.56 / 78.32 60.56 / 46.45 45.67 / 39.36 92.69 / 89.61 66.59 / 63.60 48.96 / 41.32 ICL 93.31 / 82.95 65.81 / 49.75 52.70 / 45.59 68.86 / 60.84 62.19 / 55.58 62.78 / 59.43 InComeS 91.16 / 76.81 65.15 / 45.66 55.22 / 53.97 97.22 / 87.09 70.70 / 52.23 61.33 / 59.66 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 - 36.21 / 36.21 43.86 / 43.86 - FT-M 98.93 / 90.18 49.39 / 43.13 15.73 / 12.93 99.51 / 92.60 50.04 / 46.41 18.00 / 27.27 LoRA [16] 77.31 / 72.22 37.04 / 31.65 0.380 / 02.24 86.88 / 77.78 28.61 / 24.13 1.290 / 0.330 ROME [24] 92.69 / 34-33 40.25 / 34-33 38.76 / 34-33 97.86 / 34-33 50.43 / 34-33 51.38 / 34-33 R-ROME [10] 92.59 / 34-33 40.15 / 34-33 38.70 / 34-33 97.89 / 34-33 50.47 / 34-33 51.31 / 34-33 MEMIT [25] 34-33 / 91.16 34-33 / 39.85 34-33 / 25.85 34-33 / 93.28 34-33 / 49.97 34-33 / 51.85 EMMET [11] 34-33 / 88.05 34-33 / 39.37 34-33 / 100.0 34-33 / 93.64 34-33 / 48.44 34-33 / 100.0 MEND [28] 34-33 / 35.13 34-33 / 15.29 34-33 / 07.29 34-33 / 50.91 34-33 / 38.83 34-33 / 47.12 GRACE [12] 31.34 / 33.77 25.60 / 18.55 19.19 / 07.29 33.27 / 26.79 14.35 / 11.25 12.31 / 23.45 KN [3] 36.33 / 34-33 29.60 / 34-33 32.79 / 34-33 14.49 / 34-33 8.49 / 34-33 33.21 / 34-33 IKE [48] 96.40 / 34-33 75.33 / 34-33 46.98 / 34-33 99.75 / 34-33 83.17 / 34-33 43.19 / 34-33 SERAC [29] 91.79 / 80.68 51.12 / 41.26 37.84 / 35.10 91.12 / 82.56 62.41 / 52.63 40.56 / 42.46 ICL 90.24 / 85.28 66.99 / 51.66 52.63 / 40.97 71.75 / 71.57 66.10 / 64.57 51.43 / 47.12 InComeS 90.96 / 71.44 66.69 / 47.93 52.04 / 42.02 97.95 / 91.29 75.63 / 61.22 57.16 / 59.21
@@
Table 7: More results for $WikiData_{counterfact}$ [2, 46] and ZsRE-extended [43, 46]. The data format of each cell is in "single-edit result / 100-edits result".
C.1 Datasets

### v3

@@ 4.2 Main results
Llama-3.2-1B Method Time (seconds) InComeS-Compression 0.0326 ICL-Prefilling 0.8934 FT-M 3.4124 LoRA 33.5412 MEMIT 112.2238 EMMET 158.6524 Qwen2.5-7B InComeS-Compression 0.1071 ICL-Prefilling 0.9082 FT-M 28.6132 LoRA 122.1876 MEMIT 423.4823 EMMET 512.4235
Table 4: Measured time (seconds) for 100 edits.
Method Time (seconds) Encoding w/ $1k$ InComeS-Compression 0.2108 ICL-Prefilling 1.0413 Decoding w/ $1k$ InComeS-Selection 0.0274 ICL-Generation (with prefilled cache) 0.1555 Encoding w/ $2k$ InComeS-Compression 0.4051 ICL-Prefilling 1.2165 Decoding w/ $2k$ InComeS-Selection 0.0297 ICL-Generation (with prefilled cache) 0.3545
Table 5: Scaled efficiency comparison (seconds) between InComeS and ICL using Llama-3.2-1B.
4.3 Efficiency Analysis
Finally, we present the efficiency analysis for our method. By default, the individual edit length is around 10 to 11. We first compare the efficiency of our method with the efficiency of other knowledge editing methods. Table 4 reports the time required to perform 100 edits for each method. Our method has significantly better efficiency than the other presented editing methods. Additionally, compared to ICL, our approach only needs to maintain the KV cache of the gist representations from the deeper half layers, resulting in substantially lower memory cost. To verify our method’s superiority in efficiency on long context, we further conduct experiments on scaled context length (Table 5). The result demonstrates the efficiency advantage of our method in both the encoding and decoding stages. More detailed analysis can be found in Appendix C.3.
@@ 4.4 Ablation study & Analysis
Method MQuAKE Single editing Batch editing InComeS 71.99 53.15 - w/o zero-gist 58.98 40.59 - w/ full model 72.40 52.34 - w/o loss on query 59.86 44.78 - w/ golden loss 71.09 48.56 - w/o kl loss (Eq. 2) 55.10 51.71 - w/o token weighting (Eq. 1) 55.02 52.63
Table 6: Ablation and Analysis experiments, the edit batch size is 100 for all results. Detailed results can be found in Table 8.
@@
We present the reason for our decision to use the KV cache from the second half of the model layers. To investigate this, we train a model using the KV cache from all layers and evaluate it on 1000 instances from ZsRE Levy et al. (2017). We record the probabilities allocated to the zero-gist in the cross-attention modules, as shown in Fig. 4b. The result shows that the zero-gist probabilities in layers 7-15 are generally lower than those in layers 1-6, and there is a notable drop in the zero-gist probability at layer 7. This suggests that, even when trained to use the full-model KV cache, the model mainly relies on information from deeper layers, since higher probabilities of the zero-gists indicate lower utilization of the actual edit contexts. A possible explanation is that more information is accumulated in the deeper layers, which aids both compression and selection processes. To verify our analysis, we also test the full layer trained model (see the “w/ full model” line in Table 6). The results show nearly no increase compared to the half model case (InComeS). Additionally, restricting the KV cache to only the second half of the model could provide efficiency benefits with lower memory and computation costs.
@@
Applying a small temperature to the gist cross-attention sharpens the probability distribution over the gist KV caches, which facilitates the model’s ability to retrieve the correct information. We determine the appropriate temperature based on entropy, which has been shown to be an important factor in attention mechanisms Zhang et al. (2024b). Specifically, we aim to keep the cross-attention entropy close to its optimal value, which occurs when it only needs to attend to one edit. To achieve this, we select 1000 instances from ZsRE Levy et al. (2017) and calculate the entropy of the edit batch size 1. We then calculate the entropy for larger edit batch sizes (10, 20, 40, 60, 80, 100, 200, 300) and find the temperatures that align their entropy with the optimal case via gradient descent  [^5] Note that all entropy is calculated in a way that the query and answer part are just a copy of the context. We report the calculated temperature in Figure 4a. As expected, the temperature decreases as the edit batch size increases, but interestingly, it gradually converges to a specific value. Specifically, layers 9, 10, and 14 finally converge to around 0.5. To encourage more decisive selection, we slightly lower this value and set the temperature to $T=0.45$.
@@
More analysis
More analysis including side-effect analysis, etc. is provided in Appendix C.
@@ 5 Related Work
The area of knowledge editing (or model editing) has experienced a thriving development in recent years. Researchers have explored various directions in this area. One typical direction is to adopt external memory for the edits. The memory formats applied by different researchers are diverse. Methods like SERAC Mitchell et al. (2022b), IKE Zheng et al. (2023), DR-IKE Nafee et al. (2025), MeLLo Zhong et al. (2023a) adopt explicit non-parametric memory, which stores specific edit instances, and a retriever that is responsible for recalling relevant edits from the memory. For example, IKE uses KNN, and SERAC applies a trained classifier. Another line of work, such as RECIPE Chen et al. (2024), applies implicit parametric memory to store the edits. CaliNET Dong et al. (2022), T-Patcher Huang et al. (2023) embeds the knowledge into a fixed number of neurons and adds them to the model. GRACE Hartvigsen et al. (2022) adopts a discrete key-value codebook with the value optimized for the desired knowledge. MELO Yu et al. (2024) applies dynamic LoRA blocks and indexes them via an internal vector database. KE De Cao et al. (2021), MEND Mitchell et al. (2022a) train a separate meta-model for editing. Another popular direction is to merge knowledge into the model directly. Methods like KN Dai et al. (2022), ROME Meng et al. (2022), R-ROME Gupta et al. (2024a), MEMIT Meng et al. (2023), PMET Li et al. (2023), CoachHooK Li et al. (2024), and AlphaEdit Fang et al. (2024) perform editing by tweaking the located FFN part of the model directly. However, some studies reveal that these methods could potentially bring about side effects in the original model Gu et al. (2024); Pinter and Elhadad (2023), leaving the real effectiveness of these methods to be further investigated.
@@ 6 Conclusion
In this paper, we propose InComeS, a scalable model editing method that integrates compression and selection mechanisms directly into the LLMs. InComeS adopts a context compression technique to condense the editing context to KV representations on top of the introduced gist tokens and takes advantage of the compressed KVs to efficiently retrieve the relevant editing context information. Experiments on four different and complex editing settings demonstrate the superiority of our method for comprehensively editing. Further Analysis and ablations validate each component of InComeS and demonstrate the great efficiency and performance gain of our method.
7 Limitations
@@
Due to the limited computational resources, we only extend the model size to 7B and leave the larger model size to future work. We are aware that the original gisting work Mu et al. (2023) conducts experiments on three model architectures, i.e., encoder-decoder, encoder-only, and decoder-only. In this work, we determine to focus on the decoder-only autoregressive architecture as it is the structure used by most of the popular models nowadays Yang et al. (2024); OpenAI (2023); DeepSeek-AI et al. (2024).
@@
In this work, we maintain the compression rate to roughly around 12:1 (one edit, which contains around 12 tokens for all testing datasets except DUNE Akyürek et al. (2023), corresponds to one gist token), as one edit represents a fine-grained piece of information. However, we believe it is necessary to investigate the impact of lowering the compression rate, since it potentially helps extend the length of a single edit Deng et al. (2024, 2025).
@@
Appendix A Training details
@@
InComeS is trained on around 1.5 billion tokens, which mainly come from summarization and QA datasets. Specifically, for summerization datasets, we select $4.5e^{6}$ instances from S2ORC Lo et al. (2020), $1.15e^{6}$ instances from AG News Corpus  [^6] https://huggingface.co/datasets/sentence-transformers/agnews; and for QA datasets, we use squad Rajpurkar et al. (2016), a modified version  [^7] https://huggingface.co/datasets/LLukas22/nq-simplified of the natural question dataset Kwiatkowski et al. (2019), OpenBookQA Mihaylov et al. (2018), QASC Khot et al. (2020), MedMCQA Pal et al. (2022), and NetEval Miao et al. (2023). We also include the training split of ZsRE Levy et al. (2017), COUNTERFACT Meng et al. (2023), $Wiki_{counterfact}$ from EasyEdit framework Wang et al. (2023).
@@
We use a cosine linear-warmup scheduler for both models, with a maximum learning rate $1e^{-5}$ and a minimum learning rate $1e^{-6}$ for Llama-3.2-1B and a maximum $5e^{-6}$ and a minimum $1e^{-6}$ for Qwen2.5-7B. To improve the model’s robustness and sample them at a predefined rate during training, the batch size is dynamically sampled from a predefined set rather than a fixed number. Specifically, the predefined set for batch size is 8, 16, 32, 64, and 128, and their corresponding sample rates are 0.05, 0.05, 0.05, 0.15, and 0.7. We adopt DeepSpeed Rajbhandari et al. (2020, 2021); Ren et al. (2021) and Liger Kernel Hsu et al. (2024) with 8 Nvidia GPUs for distributed training. Overall, the training takes around 11 hours for Llama-3.2-1B and 35 hours for Qwen2.5-7B.
Appendix B Experiment details
@@
Method Model $\text{WikiData}_{counterfact}$ ZsRE-extended Edit Success Portability Locality Edit Success Portability Locality Base Llama-3.2-1B 21.28 / 21.28 19.73 / 19.73 - 30.06 / 30.06 40.17 / 40.17 - FT-M 97.02 / 94.58 53.43 / 47.51 49.01 / 20.54 99.81 / 95.94 62.80 / 54.84 75.59 / 59.28 LoRA 98.91 / 82.61 52.87 / 43.84 23.31 / 14.94 99.86 / 93.18 57.43 / 44.85 34.94 / 37.41 ROME 94.33 / 44-33 40.44 / 44-33 26.13 / 44-33 95.41 / 44-33 46.04 / 44-33 40.19 / 44-33 R-ROME 94.31 / 44-33 40.65 / 44-33 25.83 / 44-33 95.29 / 44-33 46.46 / 44-33 39.95 / 44-33 MEMIT 77.87 / 66.94 46.34 / 23.51 33.45 / 11.12 79.43 / 58.79 43.26 / 25.68 40.33 / 30.10 EMMET 34.56 / 27.02 15.43 / 08.86 100.0 / 100.0 23.68 / 15.96 17.54 / 06.84 100.0 / 100.0 MEND 38.45 / 26.66 24.98 / 21.06 17.34 / 13.54 53.45 / 43.33 37.53 / 30.77 46.31 / 38.75 GRACE 33.27 / 25.06 14.33 / 10.51 28.71 / 11.43 32.00 / 24.44 12.73 / 10.91 24.12 / 10.12 KN 20.60 / 34-33 17.16 / 34-33 19.46 / 43-33 16.01 / 43-33 06.70 / 43-33 21.23 / 34-33 IKE 61.70 / 34-33 45.55 / 34-33 48.80 / 43-33 59.15 / 43-33 57.39 / 43-33 40.21 / 34-33 SERAC 89.56 / 78.32 60.56 / 46.45 45.67 / 39.36 92.69 / 89.61 66.59 / 63.60 48.96 / 41.32 ICL 93.31 / 82.95 65.81 / 44.75 52.70 / 45.59 68.86 / 60.84 62.19 / 51.58 62.78 / 59.43 InComeS 91.16 / 76.81 65.15 / 45.66 55.22 / 53.97 97.22 / 87.09 70.70 / 52.23 61.33 / 59.66 Base Qwen2.5-7B 22.35 / 22.35 21.46 / 21.46 - 36.21 / 36.21 43.86 / 43.86 - FT-M 98.93 / 90.18 49.39 / 43.13 15.73 / 12.93 99.51 / 92.60 50.04 / 46.41 18.00 / 27.27 LoRA 77.31 / 72.22 37.04 / 31.65 0.380 / 02.24 86.88 / 77.78 28.61 / 24.13 1.290 / 0.330 ROME 92.69 / 34-33 40.25 / 34-33 38.76 / 34-33 97.86 / 34-33 50.43 / 34-33 51.38 / 34-33 R-ROME 92.59 / 34-33 40.15 / 34-33 38.70 / 34-33 97.89 / 34-33 50.47 / 34-33 51.31 / 34-33 MEMIT 93.43 / 91.16 43.58 / 39.85 33.79 / 25.85 95.23 / 93.28 53.23 / 49.97 54.23 / 51.85 EMMET 92.33 / 88.05 43.24 / 39.37 100.0 / 100.0 94.32 / 93.64 51.55 / 48.44 100.0 / 100.0 MEND 46.73 / 35.13 20.45 / 15.29 13.34 / 07.29 54.32 / 50.91 43.26 / 38.83 54.39 / 47.12 GRACE 31.34 / 33.77 25.60 / 18.55 19.19 / 07.29 33.27 / 26.79 14.35 / 11.25 12.31 / 23.45 KN 36.33 / 34-33 29.60 / 34-33 32.79 / 34-33 14.49 / 34-33 8.49 / 34-33 33.21 / 34-33 IKE 96.40 / 34-33 65.33 / 34-33 46.98 / 34-33 99.75 / 34-33 70.17 / 34-33 43.19 / 34-33 SERAC 91.79 / 80.68 51.12 / 41.26 37.84 / 35.10 91.12 / 82.56 62.41 / 52.63 40.56 / 42.46 ICL 90.24 / 85.28 66.99 / 51.66 52.63 / 40.97 71.75 / 71.57 66.10 / 60.57 51.43 / 47.12 InComeS 90.96 / 71.44 66.69 / 47.93 52.04 / 42.02 97.95 / 91.29 75.63 / 61.22 57.16 / 59.21
@@
Table 7: More results for $WikiData_{counterfact}$ Cohen et al. (2024); Zhang et al. (2024a) and ZsRE-extended Yao et al. (2023); Zhang et al. (2024a). The data format of each cell is in "single-edit result / 100-edits result".
Method MQuAKE Single editing Batch editing 2-edits 3-edits 4-edits Avg 2-edits 3-edits 4-edits Avg Llama-3.2-1B InComeS 71.19 72.17 72.62 71.99 53.93 52.79 52.73 53.15 - w/o zero-gist 62.00 61.93 53.01 58.98 47.48 41.71 32.57 40.59 - w/ full model 69.74 70.02 77.43 72.40 51.91 52.59 52.53 52.34 - w/o loss on query 60.92 60.97 57.70 59.86 49.63 46.44 38.28 44.78 - w/ golden loss 69.91 69.98 73.37 71.09 55.39 50.24 40.06 48.56 - w/o kl loss (Eq. 2) 57.75 57.38 50.18 55.10 52.47 50.86 51.79 51.71 - w/o token weighting (Eq. 1) 57.45 57.84 49.76 55.02 53.49 53.20 51.91 52.63
Table 8: Full results of the ablation and analysis experiments.
Method Single Editing Batch Editing 2-edits 3-edits 4-edits Avg 2-edits 3-edits 4-edits Avg Llama-3.2-1B Base 41.79 43.51 31.58 38.96 41.79 43.51 31.58 38.96 FT-M 55.32 56.59 52.22 54.71 51.89 50.15 44.1 48.71 LoRA 67.63 68.84 60.21 65.56 50.29 47.88 47.15 48.44 ROME 1.97 7.77 0.55 3.43 - - - - R-ROME 2.72 4.29 8.53 5.18 - - - - MEMIT 40.21 39.54 22.34 34.03 29.71 28.87 15.7 24.76 EMMET 5.4 6.61 4.74 5.58 11.49 17.16 15.89 14.85 GRACE 8.45 7.13 5.24 6.94 2.03 2.68 2.06 2.26 SERAC 41.84 43.51 31.58 38.97 41.84 43.45 31.74 39.01 MEND 39.88 35.23 33.31 36.41 35.55 30.29 29.89 31.91 RECIPE 58.44 54.69 53.33 55.49 50.71 47.34 39.39 45.81 DR-IKE 52.95 48.26 47.12 49.44 42.45 41.33 32.22 38.67 ICL 59.23 59.00 51.63 56.62 50.06 49.86 42.37 47.43 InComeS 71.19 72.17 72.62 71.99 53.93 52.79 52.73 53.15 Qwen2.5-7B Base 44.08 44.14 30.62 39.61 44.08 44.14 30.62 39.61 FT-M 69.89 74.88 74.96 73.24 50.31 49.04 50.39 49.91 LoRA 36.95 34.28 29.02 33.42 16.41 24.22 23.12 21.25 ROME 9.67 7.33 8.93 8.64 - - - - R-ROME 10.73 6.68 3.62 7.01 - - - - MEMIT 44.14 46.05 32.15 40.78 43.94 46.10 30.55 40.20 EMMET 26.10 38.38 26.36 30.28 40.83 45.02 33.89 39.91 GRACE 15.78 13.45 13.23 14.15 5.43 7.88 4.23 5.85 SERAC 55.56 59.23 53.67 56.15 42.34 40.33 39.39 40.69 MEND 34.23 45.34 30.25 36.61 39.88 35.45 34.21 36.51 RECIPE 57.54 59.69 58.43 58.55 49.24 49.78 40.71 46.58 DR-IKE 60.95 54.66 51.22 55.61 44.35 39.99 39.77 41.37 ICL 69.76 76.91 74.54 73.74 53.53 50.54 44.77 49.61 InComeS 66.46 71.24 76.54 71.41 55.13 53.48 47.91 52.17
Table 9: Full results on MQuAKE Zhong et al. (2023b).
Method Single Editing Batch Editing New info Scientific R. Debiasing Avg New info Scientific R. Debiasing Avg Llama-3.2-1B Base 56.85 55.87 32.73 48.48 56.85 55.87 32.73 48.48 FT-M 57.43 53.34 35.73 48.83 57.07 53.45 33.43 47.89 LoRA 53.65 50.87 36.73 47.08 56.77 54.66 35.83 49.08 SERAC 52.67 48.96 46.73 49.45 50.76 47.32 36.52 44.87 ICL 58.67 55.84 56.65 57.05 56.94 55.68 39.46 50.69 InComeS 60.00 58.17 54.61 57.59 57.76 56.46 46.14 53.45 Qwen2.5-7B Base 63.44 66.03 36.95 55.47 63.44 66.03 36.95 55.47 FT-M 64.83 67.77 48.73 60.44 64.27 64.56 41.51 56.78 LoRA 63.85 63.22 39.56 55.54 62.58 65.33 43.73 57.21 SERAC 64.45 63.57 33.38 53.8 56.78 58.97 31.24 48.99 ICL 65.81 65.34 35.25 55.47 66.29 66.24 40.73 57.75 InComeS 66.83 68.02 62.59 65.81 65.61 67.82 56.69 63.37
Table 10: Full results on DUNE Akyürek et al. (2023).
B.1 Datasets